OpenStudy (anonymous):

If r and s are positive numbers satisfying the inequality r/s<(r+1)/(s+1), which of the following MUST be true? r1 s>1 r and s can be any number as long as r≠s

OpenStudy (anonymous):

This sounds like your everyday SAT problem. It's a logic question. Pick two random numbers and plug in and try the different conditions til one works!

OpenStudy (anonymous):

r/s<(r+1)/(s+1) \(\large \frac{r}{s}=\frac{r+1}{s+1} \rightarrow r(s+1)<s(r+1) \) rs + r < rs + s r < s